OpenStudy (anonymous):

What is the equation of a circle that passes through point J(6, 4) and has center K (1, -8).

OpenStudy (anonymous):

equation of a circle with radius \(\large r\) and the center \(\large (a,b)\) is \[\large (x-a)^2+(y-b)^2=r^2\]

OpenStudy (anonymous):

now u tell me what are the values of a and b?

OpenStudy (anonymous):

i got 169 as the answer.

OpenStudy (anonymous):

plz show your work

OpenStudy (anonymous):

u mean r^2=169 am i right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

exactly so r=13

OpenStudy (anonymous):

and the equation will be \[\large (x-1)^2+(y+8)^2=169\] am i right?

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

almost done